Let y=0.25x + 4 for the cost of going to a movie from 1985 to 1995. Let x represent the number of years since 1985. What is the y-intercept in the linear model? What is the estimated cost of going the movies in 1991? What is the estimated cost of going to the movies in 1997?
Answer by mananth(16946)

Let y=0.25x + 4 for the cost of going to a movie from 1985 to 1995. Let x represent the number of years since 1985. What is the y-intercept in the linear model? What is the estimated cost of going the movies in 1991? What is the estimated cost of going to the movies in 1997?
..
y=0.25x+4 1985 ,x=0
y intecept = 4
..
1991, x=6
y=0.25*6 +4
y=1.5+4
y=5.5
...
applying the same linear model 1997 x=12
y=0.25*12+4
y=3+4
y=7
...
m.ananth@hotmail.ca